A dinner plate fits snugly in a square box with perimeter 48 inches . what is the circumference of the plate ?
Answer by checkley77(12844)

48/4=12 is the lengtth of the sides.
12 inches is the diameter of the circle.
C=2PIR
C=2*3.14*12/2
C=6.28*6
C=37.68 INCHES ANS.